Engineering Night UT 20180401

Night Plan:

  • After Opening:
    • DONE AO Startup with OSA (Doug, Al, Steve)
    • DONE Verification of Conditions (Doug, Al, Steve)
    • DONE Total Time for AO startup activities ~ 1 hour

  • IF CLEAR & Seeing < 1.2":
    • AO Engineering Tests - HIGHEST PRIORITY
      • DONE Collected more data for TT LUT
      • w-unit transforms for AGW1 and AGW2 ~ confirmation test ~30min (Doug, Barry)
      • W-unit transforms for AGW1 and AGW2 ~ confirmation test ~30 min (Doug, Barry)
    • Field Aberration data for both sides, although DX is the primary target (John H., Barry)
      • The new pointing and collimation models from DougS should be installed prior to collecting field aberration data.

  • ANY Seeing Conditions:
    • DONE Non-Sidereal test - IT 7081 ~ 45 minutes (including overheads) - MODS-Binocular (Prepared by Al & Olga, executed by Jenny and/or Barry)
      • Before 10 UT: 617 Patroclus
      • After 10 UT: JVII Elara

  • Poor Seeing and/or significant clouds (extinction ~ few mags)
    • Collimation and Pointing models -MODS-Binocular-Blue ~ 1-2 hours HIGHEST PRIORITY FOR POOR CONDITIONS
      • DONE The new pointing and collimation models from DougS should be installed prior to collecting more pointing data.
    • TMS Testing - LBC (J. Hill, A. Rakich) ~ any amount of time
    • WRS Testing - LBC (O. Kuhn) ~ any amount of time
    • Poor weather backup programs if TMS and Collimation/Pointing are completed (UT April 1?)

Night

OSA: S. Allanson

ISA:B. Rothberg

ISp: E. Solheid

AO: D. Miller

SO: J. Power

LUCI-AO Pre-check

UT 00:35 - Startng LUCI GUIs, taking some test darks

UT 00:37 - waking up MODS-1/MODS-2 Blue only

UT 00:52 - MODS sievesnaps taken

UT 01:01 - had to update the Slitmask.tab file on both MODS computers (mask exchange was done but previous slits were still listed), running maskSnap on blue only for new masks to verift they are OK.

UT 01:14 - AO Startup revisions proceeding, MODS masks exchange is now complete

UT 01:29 - Enclosure open - on-sky AO test/startup proceeding

UT 02:56 - we see but 2 stars in 'ye old skycam

LUCI-AO data collection

UT 03:24 - attempting to work on AO - N30fieldstops aligned on both LUCIs. Can't see many stars in all-sky but Doug is fairly confident he has some tests he can run

UT 03:34 - beginning collection on LUCI-2, will optimize LUCI-1 further. - LUCI-2 data is flat-topped

UT 03:49 - finding a fainter star

UT 03:53 - clouds have thickened, so staying put, data looks good

UT 04:25 - had upwards of 7mags extinction, back to about 4mags extinction

UT 05:00 - 7+mags of extinction, no stars on All-Sky no flux on stages or AO CCDs. Holding now....

UT 05:51 - sucker hole has vanished, back to pea soup

UT 06:30 - completeing AO data collection - will reconfigure shortly for MODS - Binocular for NS testing and then collimation/pointing (if we can see stars) - LUCI in safe state

MODS-NS test

UT 06:37 - Reconfiguring for MODS-Binocular

UT 07:00 Taking first images on Eugenia MOD1

The script contained offsets. Guiding did not pause on first pointing. but it did on the second.

UT 07:07 resending script to see if the missed pause on first offset (which is before any exposure) was a fluke or not.

So the lack of pause was a fluke. There was a pause for both sides on each exposure each time. Tunning scripts which is a 3x3 grid of 10" . r on the left, and g on the right. Guiding resuming too early after offsets. Several images look jumped. Performing asyns offsets and it seem slike one finidshing before the other throws things off.

UT 07:20 The guide probe is definitely not on axis so partially sucess.

UT 07:33 - seeing jumps in exposures, could be the dithering on each side is interfering, or another problem. Three troubleshooting steps: 1) keep binouclar, but no dithering; 2) switch to monocular, then do dither; and 3) monocular no dither

UT 07:34 - mods2b.20180401.0012 shows jump, mods2b.20180401.0016 shows jump, mods1r.20180401.0008 jump

UT 07:37 - resending with binocular no dithers, we did see a large jump before resending this, will take a number of exposures to see if jumps crop up

UT 07:40 - IMCS lock in progress, but guider on both sides show huge jumps - IT 6746 - these jumps occurred while we still had flux in the guiders to see the gstars - mods2b.20180401.0019, mods1r.20180401.0010 - reopening IT 6746

UT 07:41 - flux is dropping out due to thick clouds

UT 07:45 - calling Petr

UT 07:47 - MODS2 guideprobe has lost the star - unable to continue with MODS-2, continuing with MODS-1

UT 07:53 - MODS1 gstar returns, we are prepping a LUCI script to see if same behavior occurs RM507/Service_Observing/NSGuiding

UT 08:03 - Downtime, can't guide on an r=11 guidestar, Petr investigating NS jumps

UT 08:07 - Closing enclosure - we can't guide on an 11th mag star and the pointing stars are 10th mag....we'll monitor and re-open if it improves.

MODS Pointing/Collimation

PtStars output for LST 14:15

Active Preset to each, take image with MODS

UT 08:59 - moving to first target to acquire

UT 09:01 - homing both probes after NS issues

- will take a 10sec test exposure to gauge needed exposure

- need to collimate for each one - but need to make sure clouds don't obscure too much

- Update 20180401_SX.log, _DX.log,

- running from /lbt/data/config/tcs/PSF

-rwxr-xr-x. 1 telescope 1513 288 Mar 28 13:42 mar28.sh
-rwxr-xr-x. 1 telescope 1513 288 Mar 28 13:41 mar28_revert.sh
-rw-rw-r--. 1 telescope 1513 1428 Mar 28 13:37 DXSMLUCIFERCollimation.20180328.dat
-rw-rw-r--. 1 telescope 1513 1902 Mar 28 13:35 DXPMLUCIFERCollimation.20180328.dat
-rw-rw-r--. 1 telescope 1513 1383 Mar 28 13:28 SXSMLUCIFERCollimation.20180328.dat
-rw-rw-r--. 1 telescope 1513 1869 Mar 28 13:22 SXPMLUCIFERCollimation.20180328.dat
files renamed from LUCIFER to MODS

-also update the mar28.sh and mar28_revert.sh to reflect MODS instead of LUCI

-next step is load the updated pointing models /lbt/data/config/tcs/PCS --> need to update PCSInstrument.conf with ?X_Gregorrian_20180328.ptmod

UT 09:19 -restart PCS, reinitialize instrument, restarting Pointing Model, reinitialize all 6 mirrors, run mar28.sh script

manually configure MODS-1 and MODS-2 blue imaging only 3kx3k - make sure to run IMCS lock before each exposure

UT 09:22 - starting log

UT 09:23 - no star seen SX, unable to abort preset

UT 09:28 - unable to see anything on SX again, homing probe on MODS-1 to make sure

UT 09:29 - spiraling to look for SX star

UT 09:31 - spiraled to star on SX, but not sure if its the right one, moving to 2nd star to confirm

UT 09:33 - trying to PT10m068 guidestar 8

AZ ALT RA J2000 DEC PtStarID mods1b.20180401.XXXX mods2b.20180401.XXXX
======= ===================== =====================================
5.82 47.21 15:46:01.267 74:50:04.07 PT10m068 0005 0042

15.84 74.38 15:10:01.667 47:35:14.39 PT10m065 skipping no good guidestars bright enough to be seen through muck

29.49 41.95 18:44:23.495 64:56:46.84 PT10m082 0006 0043

46.93 59.07 17:05:49.687 49:21:32.03 PT10m074 0007 0044 - don't use

UT 09:59 - active optics cleared - MODS-1 image has a tail

UT 10:00 - trying new guidestar same target, new WFS image looks good,

46.93 59.07 17:05:49.687 49:21:32.03 PT10m074 0008 0045

UT 10:03 - will go back to guidestar 15 - looks better now

52.35 34.70 19:34:12.992 46:54:10.91 PT10m086 - pointing off on SX, can't even collimate

UT 10:11 - going to revert back to previous model

UT 10:16 - need to revert to last MODS collimation model - where is the last collimation model? - should have been a version number/date put in for the last version of the MODS collimation file added to the PSF directory or the shell script should have created a backup first before overwriting. BR should have verified the shell script to make sure backups were made first and/or that there was a backup in the directory to start with before making changes.

checking SVN to recover collimation model - success - need an easier recovery method from SVN for text files

UT 10:29 - Moon barely visible

UT 10:38 - restarting everything, will run a test on a target (Vega) - found - can collimate OK - we are recovered

UT 11:17 - closing up, this muck is thick, can't even see Vega clearly on the all-sky

UT 11:18 - putting both MODS to sleep and powering off LBCs

-- %USERSIG{BarryRothberg - 2018-03-31}%
Topic revision: r5 - 01 Apr 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback